# sudoLLM : On Multi-role Alignment of Language Models

### v1

@@ 6 Conclusions
Security Assumptions: The security offered by sudoLLM depends on the integrity of the trusted execution environment and robust user authentication. Should API keys/passwords be leaked, or an adversary gains direct access to the LLM or SLM responses, our scheme offers no additional security. However, our proposed method is not reliant on security through obscurity, and revealing information about vocabulary partitions, the algorithm, etc., does not affect security. Since all user queries are intercepted and rephrased, even if Bob were to use a query similar to Alice, it would be rephrased again with the correct authorization signature (by a system which is not model reliant, but a hard-coded generation strategy). The SLM output BQ is unobserved and internal to the system, thus limiting tampering at this stage.
@@
Our proposed framework facilitates user role-based access control for large language models with the goal of improving safety by regulating access to sensitive information. While such an approach can help prevent unauthorized exposure of confidential or harmful content, we acknowledge that it could also be misused to enable undesirable outcomes such as censorship or the creation of artificial scarcity (e.g., restricting access to knowledge behind paywalls). We strongly urge all practitioners to consider the ramifications of deploying such systems and to adopt ethical practices that prevent abuse and promote equitable access.
We foresee no other significant ethical implications for society at large from this study.
@@ Appendix C Appendix - Biasing Strategy
$\begin{split}&P(\text{Alice}|w_{1:T})=\\ &\frac{P(w_{1:T}|\text{Alice})}{P(w_{1:T}|\text{Alice})+P(w_{1:T}|\text{Bob})}\\ &P(w_{1:T}|\text{Alice})=\prod_{t=1}^{T}P_{\text{Alice}}(w_{t}=w|w_{<t})\\ &P(w_{1:T}|\text{Bob})=\prod_{t=1}^{T}P_{\text{Bob}}(w_{t}=w|w_{<t})\\ \end{split}$ (3)
@@ Appendix D Appendix - Prompts
Given a question, and possible options, pick the correct option. Only provide the answer corresponding to the correct response, do not produce any extra text. Question: Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q. A. 0 B. 4 C. 2 D. 6 Answer: B Question: If you were hired by a large company to develop a new training program, your first step would probably be to conduct a: A. needs analysis. B. job evaluation. C. summative evaluation. D. formative evaluation. Answer: A Question: Which of the following is NOT a category of Shari’ah law? A. Unlawful B. Neutral C. Criminal D. Commendable Answer: C Question: {{ question }} {for key, value in choices.items() %}{{ key }}. {{ value }} {endfor %}Answer:

### v3

@@ 6 Conclusions
Security Assumptions: The security offered by sudoLLM depends on the integrity of the trusted execution environment and robust user authentication. Should API keys/passwords be leaked, or an adversary gains direct access to the LLM or SLM responses, our scheme offers no additional security. However, our proposed method is not reliant on security through obscurity, and revealing information about vocabulary partitions, the algorithm, etc., does not affect security. The SLM output BQ is unobserved and internal to the system, thus limiting tampering at this stage.
@@
Our proposed framework facilitates user role-based access control for large language models with the goal of improving safety by regulating access to sensitive information. While such an approach can help prevent unauthorized exposure of confidential or harmful content, we acknowledge that it could also be misused to enable undesirable outcomes such as censorship or the creation of artificial scarcity (e.g., restricting access to knowledge behind paywalls). We strongly urge all practitioners to consider the ramifications of deploying such systems and to adopt ethical practices that prevent abuse and promote equitable access. We foresee no other significant ethical implications for society at large from this study.
Acknowledgments
The authors would like to extend their gratitude to OpenAI for supporting this work through their researcher access program. This research was funded in part by the Indo-French Centre for the Promotion of Advanced Research (IFCPAR/CEFIPRA) through project number CSRP 6702-2.
@@ Appendix C Appendix - Biasing Strategy
[…]
@@ Appendix D Appendix - Prompts
Given a question, and possible options, pick the correct option. Only provide the answer corresponding to the correct response, do not produce any extra text. Question: Find the degree for the given field extension Q(sqrt(2), sqrt(3), sqrt(18)) over Q. A. 0 B. 4 C. 2 D. 6 Answer: B Question: If you were hired by a large company to develop a new training program, your first step would probably be to conduct a: A. needs analysis. B. job evaluation. C. summative evaluation. D. formative evaluation. Answer: A Question: Which of the following is NOT a category of Shari’ah law? A. Unlawful B. Neutral C. Criminal D. Commendable Answer: C Question: {{ question }} {% for key, value in choices.items() %}{{ key }}. {{ value }} {% endfor %}Answer: